OpenStudy (anonymous):

Find the minimum distance from the point (3,0) to y= square root of x

OpenStudy (anonymous):

this one easier. any point on the graph will look like \[(x,\sqrt{x})\] and working with the square of the distance we see that you want to minimize \[(x-3)^2+(\sqrt{x}-0)^2\] or simply \[(x-3)^2+x\]

OpenStudy (anonymous):

don't really even need calc for this. you have \[f(x)=(x-3)^2+x=x^2-5x+9\] minimum value at the vertex, or else take the derivative, set it equal zero, etc

OpenStudy (anonymous):

okay thanks again

OpenStudy (anonymous):

yw